OpenStudy (anonymous):

The sum of one-third of a number, one-fourth of a number and one-ninth of the number is 50. Find the number. Please help and show how to solve.

OpenStudy (anonymous):

one-third of a number = 1/3x one-fourth of a number = 1/4x one-ninth of a number = 1/9x The SUM of the numbers = 1/3x +1/4x +1/9x = 50 \[\frac{ 1 }{ 3 }x + \frac{ 1 }{ 4 } x + \frac{ 1 }{ 9 }x = 50\] Solve for X by getting a common denominator first.